OpenStudy (anonymous):

.65 converted to percent is?

OpenStudy (anonymous):

65%

OpenStudy (jiteshmeghwal9):



OpenStudy (anonymous):

65/100 *100= 65

OpenStudy (anonymous):

nice answer.., :)

OpenStudy (anonymous):

thanks .

OpenStudy (anonymous):

@stephielee Glad to see you're learning from this experience. This is your fourth post looking for answers. Sadly, some people are just giving them to you, so I guess I can't blame you for continuing to do it.

OpenStudy (jiteshmeghwal9):

To convert a percent to decimal \(\Huge{\frac{1}{100}=0.01}\) To convert a decimal to percent \(\Huge{0.01=\frac{1}{100}}\) or \(\Huge{\frac{1}{100}\times 100=1}\)

OpenStudy (anonymous):

really im just double checking, ive already done my answers im just seeing how many i missed.

OpenStudy (anonymous):

Uh huh.

OpenStudy (anonymous):

okayyyyyy